OpenStudy (anonymous):

Help me?? The matrix C=[1,-2_-3,7] was used to encode a phrase to [7,-28,-25,-35,-2_-21,107,90,123,17] Find C^-1 and use it to decode the matrix.

OpenStudy (phi):

First find the inverse see http://www.mathsisfun.com/algebra/matrix-inverse.html and scroll down to the example for a 2 x 2 matrix. can you find the inverse of matrix C ?

OpenStudy (anonymous):

C= [-7,2_3,-1] ? @phi

OpenStudy (phi):

no. look at the example more carefully

OpenStudy (phi):

but you are on the right track...

OpenStudy (anonymous):

c= [7,2_3,1]?

OpenStudy (phi):

yes. Did you figured out the determinant ?

OpenStudy (anonymous):

1?

OpenStudy (phi):

yes. and 1 over the determinant is still 1. Most times, the determinant is not 1 and you get a fraction out front. But in this case the inverse of C is a nice matrix with no fractions. Next step is multiply Cinverse * Matrix can you do that ?

OpenStudy (anonymous):

Im confused what is the cinverse?

OpenStudy (phi):

my way of saying the inverse of C. It is the matrix you just found Cinverse= [7,2_3,1] the original matrix C= [1,-2_-3,7] is different.

OpenStudy (anonymous):

Ooh ok sorry so c= [7,-4_-9,7] ?

OpenStudy (phi):

? the question is The matrix C=[1,-2_-3,7] Find C^-1 you found C^-1 = [7 2 _ 3 1 ] now you multiply C^-1 times [7,-28,-25,-35,-2_-21,107,90,123,17]

OpenStudy (anonymous):

Oooh sorry I thought u meant something else How do i do that because there are only four numbers and more at the bottom do I keep multiplying by the same numbers orrrrr?

OpenStudy (phi):

matrices are really rows and columns so the problem is \[\left[\begin{matrix}7 & 2 \\ 3 & 1\end{matrix}\right]\left[\begin{matrix}7 & -28 &-25 & -35 & -2 \\ -21 & 107 & 90 & 123&17\end{matrix}\right]\] you have a 2 x 2 x 2 x 5 (# of rows x # columns x # of rows x # of columns) notice that the 2 x 2 in the middle are =. if they were not , you could not do the multiply because the matrices would be the wrong shape notice that first and last 2 x 5 will give the shape of the answer matrix

OpenStudy (anonymous):

this is confusing lol

OpenStudy (phi):

the answer will be 10 numbers, 5 in the top row, 5 in the bottom row that means this will be 10 separate problems. to get the 5 numbers in the top row, you multiply 7 2 (the first row of C^-1 ) times each column in the big matrix. for example, the first # will be <7 2> * < 7 -21 > you multiply corresponding numbers, then add: 7*7 + 2*-21 = 49-42= 7 7 will be the first number in the top row of the answer. the second # in the top row will be <7 2> * < -28 107 >

OpenStudy (phi):

Yes it is complicated. That is why I posted the video. You can keep watching it until it sinks in.

OpenStudy (anonymous):

Lol my brain is hurting. I know how to do some of it.. but still a little confused

OpenStudy (phi):

the second # in the top row will be <7 2> * < -28 107 > can you find this # ?

OpenStudy (phi):

multiply corresponding numbers then add

OpenStudy (phi):

corresponding numbers means first times first and second number times 2nd number.

OpenStudy (anonymous):

14 * -2996 = -41944

OpenStudy (anonymous):

I have no idea if thats correct or not lol

OpenStudy (phi):

no, you should multiply the first number from < 7 2> times the first number from < -28 107 >

OpenStudy (anonymous):

ooooh sorry -196 * 214?

OpenStudy (phi):

This is a "coded message" meaning the numbers will be between 1 (stands for A) and 26 (for Z) if you get a number outside that range (except for 0 which means blank space) it is wrong.

OpenStudy (phi):

***ooooh sorry -196 * 214?*** except once you get the two products, you add them: -196 + 214

OpenStudy (anonymous):

18?

OpenStudy (phi):

yes, 18. so far we have in the top row 7 18 to find the 3rd number in the top row, we multiply the top row of C^-1 times the 3rd column of the big matrix. that means < 7 2 > * ??? can you pick out the 3rd column ?

OpenStudy (anonymous):

<-25 90> ?

OpenStudy (phi):

yes, now can you find < 7 2 > * <-25 90>

OpenStudy (anonymous):

5?

OpenStudy (phi):

yes now what's the next number ? what two pairs of numbers do we use ?

OpenStudy (anonymous):

<7 2> * < -35 123> or do we use <3 1>?

OpenStudy (phi):

it is <7 2> * < -35 123> to help remember what to multiply you can arrange the problem like this  if that makes sense. you only use <3 1> for the bottom row of the answers.